# Measuring quantum relative entropy with finite-size effect

### v2

@@ I Introduction
The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameters. This general problem setting has been studied in various researches [14, 15, 16, 17, 18, 19]. The reference [19, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [20, 21] while it did not show its attainability. Under a general setting, the reference [16, Section 4.3] showed that the optimal performance of the estimation precision is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case. The meaning of the Cramér-Rao type bound $C$ is the following; when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$. When we have no nuisance parameter, the recent paper [22] showed the existence of an estimator whose mean square error (MSE) equals $\frac{C}{n}$ without additional term under special models. However, such an estimator does not exist in general.
@@
In this paper, as the first step, using the results for estimation with nuisance parameters in [16, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [19]. In practice, it is necessary to find an estimator to achieve the MSE $\frac{C}{n}+o(\frac{1}{n})$. In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound $\frac{C}{n}$. Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems. However, this method requires an adaptive choice of measurement. One typical method for this approach is a two-step estimator, which is composed of the pre-estimation stage and the local optimal estimation. When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [23, 15]. Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{C}{n}$. From the practical viewpoint, it is necessary to upper bound this difference by using the number $n$ of unknown copies. Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{C}{n}$. The papers [24, 25] also proposed different adaptive methods for parameter estimation, but their methods have the same problem. In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
@@
In summary, considering the finite size effect of our estimation, we need to find an estimator of the quantum relative entropy $D(\rho\|\sigma)$ to satisfy the following conditions. It asymptotically attains the Cramér-Rao type bound $V(\rho\|\sigma)$ in the sense of the order $\frac{1}{n}$ when the dimension $d$ is fixed. The difference between its mean square error and the Cramér-Rao type bound is evaluated with the number $n$ of copies. Further, the estimator works well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
@@
As the second step of this paper, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [26]. Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$. In addition, we evaluate the tail probability.
@@
As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases. Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy nor the relative varentropy is not bounded even with fixed dimension $d$. Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$. We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range. Our method extensively improves the sample complexity over the simple application of the full tomography [27]. In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis. We also show that our time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits d)$.
@@
When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7]. Our result contains the case studied in the papers [6, 7]. In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indecies. Although this case improves the sample complexity over the simple application of the full tomography, the improvement is incremental. The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction. In this special case, we show that our measurement can be restricted to Schur sampling.
@@
The remaining part of this paper is organized as follows. First, Section II derives the Cramér-Rao type bound for the estimation of quantum relative entropy. Section III presents our estimator for quantum relative entropy, and its performance. Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed. Section IV studies the sample complexity of our method when $d$ increases. Section V studies the time complexity of our method when $d$ increases. Appendices are devoted to the proofs of the results presented in Section III.
II Cramér-Rao type bound
@@
To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MES under the local unbiasedness condition. The reference [19, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$. That is, they showed that
@@
In the following, we show the equality in (1). For this aim, we review the Cramér-Rao type bound $C$ for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows. We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$. That is, our parameter needs to satisfy the following conditions.
@@
for $j=2,\ldots,d^{2}-1$. Then, we choose our parametrization $\rho_{\theta}$ as
@@
$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}\rho_{*}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
@@
Here, $(b)$ and $(c)$ follows from (7) and (8) and the relation $\langle L^{1},L^{1}\rangle_{\rho_{*}}=V(\rho_{*}\|\sigma)$, respectively. In addition, $(a)$ can be shown as follows.
$\mathrm{Tr}\rho_{\theta_{0}}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\stackrel{{\scriptstyle(a)}}{{=}}\mathrm{Tr}\int_{0}^{1}\rho_{\theta_{0}}^{s}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\rho_{\theta_{0}}^{1-s}ds$
@@
where $(a)$ follows from [28, Eq. (6.32)].
@@
$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=\mathrm{Tr}X_{j}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{j}=0.$ (13)
@@
which implies the condition (6). Therefore, since all the conditions (2) – (6) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)^{-1}$ at $\theta=\theta_{0}$. That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$. That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$. Thus, the quality in (1) holds.
III Estimator based on Schur transform
Similar to the papers [29, 3, 8, 9, 10, 11, 12, 13, 4, 30, 31], we employ Schur duality of ${\cal H}^{\otimes n}$. For this aim, we prepare several notations. A sequence of monotone-increasing non-negative integers $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{d})$ is called Young index. Although many references [31, 33, 32] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience. We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [31, Section 6.2], we have
@@ III Estimator based on Schur transform
$\rho^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (16)
$\sigma^{\otimes n}$ $=\bigoplus_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (17)
where $\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}$ is the completely mixed state on ${\cal V}_{\bm{\lambda}}$. We choose a basis $\{|u_{\bm{\lambda},j}\rangle\}_{j=1}^{d_{\bm{\lambda}}}$ of ${\cal U}_{\bm{\lambda}}$ such that $|u_{\bm{\lambda},j}\rangle$ is an eigenvector of $\sigma_{\bm{\lambda}}$.
@@
To estimate the quantum relative entropy $D(\rho\|\sigma)$, we employ the estimate $x_{\bm{\lambda},j}$ defined as $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$. This measurement can be done by Schur transform efficiently [26].
To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n,d}$. Since as shown in [31, (6.16) and (6.18)], we have
@@
holds. The mean square error of this estimator is evaluated as follows.
@@
This theorem is proved in Appendix A. The difference between the upper bound (22) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n,d}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$. Using (21), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$. $n$ times of these terms go to zero.
@@
$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (27)
$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (28)
@@
IV Analysis with increasing $d$

### v4

@@ 1 Introduction
In fact, when we intend to prepare a specific state $\sigma$, it is crucial to identify the difference between the known target state $\sigma$ and the unknown actual real state $\rho$. In this situation, we often focus on the fidelity $F(\rho,\sigma):=\mathrm{Tr}|\sqrt{\rho}\sqrt{\sigma}|$ for this aim. Many papers [14, 15, 16, 17, 18] studied this problem when $\sigma$ is a pure state. However, no paper focuses on this topic when $\sigma$ is a mixed state. Instead of the fidelity, one can evaluate the difference between $\sigma$ and $\rho$ by using the quantum relative entropy $D(\rho\|\sigma)$ because of the bounding inequality $-2\log F(\rho,\sigma)\leq D(\rho\|\sigma)$ [19, Eq.(3.49)]. That is, our problem works for evaluating this difference.
@@
The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameters. This general problem setting has been studied in various researches [20, 21, 22, 23, 24, 25]. The reference [25, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [26, 27] while it did not show its attainability. Under a general setting, the reference [22, Section 4.3] showed that the optimal precision of the estimation is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case. The meaning of the Cramér-Rao type bound $C$ is the following: when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$. When we have no nuisance parameter, the recent paper [28] showed the existence of an estimator whose mean square error (MSE) equals $\frac{C}{n}$ without additional term under special models. However, such an estimator does not exist in general.
@@
In this paper, as the first step, using the results for estimation with nuisance parameters in [22, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [25]. In practice, it is necessary to find an estimator to achieve the MSE $\frac{C}{n}+o(\frac{1}{n})$. In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound $\frac{C}{n}$. Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems. However, this method requires an adaptive choice of measurement. A typical method is the two-step estimator, consisting of a pre-estimation stage and a local optimal estimation stage. When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [29, 21]. Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{C}{n}$. From the practical viewpoint, it is necessary to upper bound this difference by using the number $n$ of unknown copies. Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{C}{n}$. The papers [30, 31] also proposed different adaptive methods for parameter estimation, but their methods have the same problem. In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
@@
In summary, to account for finite-size effects in our estimation, we need to find an estimator of the quantum relative entropy $D(\rho\|\sigma)$ to satisfy the following conditions. It asymptotically attains the Cramér-Rao type bound $V(\rho\|\sigma)$ in the sense of the order $\frac{1}{n}$ when the dimension $d$ is fixed. The difference between its mean square error and the Cramér-Rao type bound is evaluated in terms of the number $n$ of copies. Furthermore, the estimator performs well even when the dimension $d$ of the quantum system increases in addition to the number $n$ of the prepared copies.
@@
As the second step of this paper, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [32]. Then, we evaluate its mean square error and compare it to the Cramér-Rao type bound with a finite number $n$. In addition, we evaluate the tail probability.
@@
As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases. Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy or the relative varentropy is not bounded even with fixed dimension $d$. Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$. We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range. Our method extensively improves the sample complexity over the simple application of the full tomography [33]. In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis. We also show that our time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits d)$.
@@
When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7]. Our result contains the case studied in the papers [6, 7]. In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indices. Although this case improves the sample complexity over the simple application of the full tomography, the improvement is incremental. The paper [7] also showed a lower bound $O(d)$ for the sample complexity, without restricting our algorithm to EYD algorithm. In this special case, we show that our measurement can be restricted to Schur sampling.
@@
The remaining part of this paper is organized as follows. First, Section 2 derives the Cramér-Rao type bound for the estimation of quantum relative entropy. Section 3 presents our estimator for quantum relative entropy, and its performance. Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed. Section 4 studies the sample complexity of our method when $d$ increases. Section 5 studies the time complexity of our method when $d$ increases. Appendices are devoted to the proofs of the results presented in Section 3.
2 Cramér-Rao type bound
@@
To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MSE under the local unbiasedness condition. The reference [25, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$. That is, they showed that
@@
In the following, we show the equality in (1). For this aim, we review the Cramér-Rao type bound $C$ for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows. We choose a parameterization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$. That is, our parameter needs to satisfy the following conditions.
@@
for $j=2,\ldots,d^{2}-1$. Then, we choose our parameterization $\rho_{\theta}$ as
@@
$=$ $\mathrm{Tr}\frac{\partial\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}(\log\rho_{*}-\log\sigma)$
$+\mathrm{Tr}\rho_{*}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
@@
Here, $(b)$ and $(c)$ follow from (7) and the pair of (8) and the relation $\langle L^{1},L^{1}\rangle_{\rho_{*}}=V(\rho_{*}\|\sigma)$, respectively. In addition, $(a)$ can be shown as follows.
$\mathrm{Tr}\rho_{\theta_{0}}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\mathrm{Tr}\int_{0}^{1}\rho_{\theta_{0}}^{s}\frac{\partial\log\rho_{\theta}}{\partial\theta^{1}}\Big{|}_{\theta=\theta_{0}}\rho_{\theta_{0}}^{1-s}ds$
@@
where $(a)$ follows from [19, Eq. (6.32)].
@@
$\frac{\partial D(\rho_{\theta}\|\sigma)}{\partial\theta^{j}}\Big{|}_{\theta=\theta_{0}}=$ $\mathrm{Tr}X_{j}(\log\rho_{*}-\log\sigma)+\mathrm{Tr}X_{j}$
$=$ $0$ (13)
@@
which implies the condition (6). Therefore, since all the conditions (2) – (6) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)$ at $\theta=\theta_{0}$. That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$. That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$. Thus, the quality in (1) holds.
3 Estimator based on Schur transform
Similar to the papers [34, 3, 8, 9, 10, 11, 12, 13, 5, 35, 36], we employ Schur duality of ${\cal H}^{\otimes n}$. For this aim, we prepare several notations. A sequence of monotone-increasing non-negative integers $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{d})$ is called Young index. Although many references [36, 37, 38] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience. We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [36, Section 6.2], we have
@@ 3 Estimator based on Schur transform
$\rho^{\otimes n}$ $=\sum_{\bm{\lambda}\in Y_{d}^{n}}\rho_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (16)
$\sigma^{\otimes n}$ $=\sum_{\bm{\lambda}\in Y_{d}^{n}}\sigma_{\bm{\lambda}}\otimes\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits},$ (17)
where $\rho_{\bm{\lambda},\mathop{\rm mix}\nolimits}$ is the completely mixed state on ${\cal V}_{\bm{\lambda}}$. In the above equations, we consider that matrices on the subspace ${\cal U}_{\bm{\lambda}}\otimes{\cal V}_{\bm{\lambda}}$ are matrices on the whole space ${\cal H}^{\otimes n}$. In the first step, we apply the Schur transform $U_{Schur}$, which realizes a basis change to block-diagonalize $U^{\otimes n}$ for an arbitrary unitary $U$. It is known that the Schur transform $U_{Schur}$ can be efficiently done [32]. We choose a basis $\{|u_{\bm{\lambda},j}\rangle\}_{j=1}^{d_{\bm{\lambda}}}$ of ${\cal U}_{\bm{\lambda}}$ such that $|u_{\bm{\lambda},j}\rangle$ is an eigenvector of $\sigma_{\bm{\lambda}}$.
@@
To estimate the quantum relative entropy $D(\rho\|\sigma)$, we employ the estimate $x_{\bm{\lambda},j}$ defined as $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$. Our estimation protocol is summarized as Protocol 1.
Protocol 1 Estimation of relative entropy
  1. Step 1.
    We apply the Schur transform $U_{Schur}$.
  2. Step 2.
    We measure the system by using the measurement $M=\{M_{\bm{\lambda},j}\}_{\bm{\lambda},j}$.
  3. Step 3.
    We set our estimate to be $x_{\bm{\lambda},j}:=-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$.
To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n,d}$. Since as shown in [36, (6.16) and (6.18)], we have
@@
holds.
The reason of our choices of the POVM $M_{\bm{\lambda},j}$ and the estimate $x_{\bm{\lambda},j}$ is the following. The reference [34] essentially showed the relation
$D(\rho^{\otimes n}\|\sigma^{\otimes n})-\frac{d(d-1)}{2}\log(n+1)$
$\leq$ $D_{M}(\rho^{\otimes n}\|\sigma^{\otimes n})\leq D(\rho^{\otimes n}\|\sigma^{\otimes n}),$ (22)
where
$D_{M}(\rho^{\otimes n}\|\sigma^{\otimes n})$
::=:= $\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}(\log\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}-\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}).$
This relation shows that the POVM $M_{\bm{\lambda},j}$ almost preserves the information with respect to the relative entropy $D(\rho^{\otimes n}\|\sigma^{\otimes n})$. Further, since the Shannon entropy $H(P_{\rho,\sigma}^{(n)}(\bm{\lambda},j)):=-\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}\log\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}$ satisfies
$0\leq H(P_{\rho,\sigma}^{(n)}(\bm{\lambda},j))\leq\log d_{n,d},$ (23)
the the combination of (21) and (22) implies
$D(\rho^{\otimes n}\|\sigma^{\otimes n})-(d+1)(d-1)\log(n+1)$
$\leq$ $-\sum_{\bm{\lambda},j}\mathrm{Tr}M_{\bm{\lambda},j}\rho^{\otimes n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$
$\leq$ $D(\rho^{\otimes n}\|\sigma^{\otimes n}).$ (24)
That is, the average of $-\frac{1}{n}\log\mathrm{Tr}M_{\bm{\lambda},j}\sigma^{\otimes n}$ is close to the value $D(\rho\|\sigma)$ to be estimated. This is the reason why we choose the estimate $x_{\bm{\lambda},j}$.
The mean square error of this estimator is evaluated as follows.
@@
This theorem is proved in Appendix A. The difference between the upper bound (25) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}(\log d_{n,d})^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$. Using (21), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$. $n$ times of these terms go to zero.
@@
$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $(d_{n,d})^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (30)
$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}$
$+d_{n,d}e^{-nr}.$ (31)
@@
4 Analysis with increasing $d$